Use any method to solve.
x/2 - y/4 = -4
x + y = -2
The answer is (-6, 4), how do I get to this answer.

Found 2 solutions by richwmiller, stanbon:
Answer by richwmiller(17219)

multiply the first by 4
2x-y=-16
x + y = -2
add them
3x=-18
x=-6
x + y = -2
-6 + y = -2
y=4

Answer by stanbon(75887)

x/2 - y/4 = -4
x + y = -2
The answer is (-6, 4), how do I get to this answer.
-------
Multiply thru the top equation by 4 to get:
2x - y = -16
x + y = -2
-------
Add and solve for "x":
3x = -18
x = -6
------
Solve for "y":
x + y = -2
-6 + y = -2
y = 4
---------------------
Solution: (-6,4)
Cheers,
Stan H.
====================